### v1

@@ 1 Introduction
  • Empirical validation. To validate our theoretical observations, in Section 6, we conduct extensive numerical experiments, leveraging sparsification operators such as Top-$K$ and Sparsified-$K$. These experiments corroborate the effectiveness of CompFedRL in practice, further reinforcing the significance of our contributions in advancing communication-efficient FedRL algorithms. Our experiments’ code is anonymously available at https://github.com/AliBeikmohammadi/CompFedRL.
@@ 2 Related Work
Single-agent RL Algorithms.
@@
Distributed and Federated RL Algorithms.
@@
Communication-Efficient Learning Algorithms.
Many compression schemes, i.e. sparsification and/or quantization, have been used for optimization algorithms to train state-of-the-art learning models (e.g. AlexNet, VGG19, and ResNet152) under resource-constrained networks [1, 64, 61, 41]. These compression schemes can be either unbiased (such as QSGD [1], TernGrad [64], and ATOMO [61]) or biased (e.g. Top-$K$[2], normalized gradient [40] and sign compression [9]). Despite communication benefits, optimization algorithms with compression often have poor convergence performance. For instance, the biased compression algorithms are shown to fail to converge for simple problems [11, 33]. To address the problem of direct compression, several studies [51, 33, 2, 32, 38, 28, 43] have proposed error-feedback mechanisms which leverage memory of previous compression errors to mitigate compression biases for distributed and federated learning algorithms under various settings. None of the mentioned works has investigated communication-efficient learning in the context of RL. The only existing study, [44], examines error-feedback compression, focusing on TD learning with linear function approximation in distributed learning, contrasting with our synchronous federated $Q$-learning approach, where aggregation occurs periodically. We investigate both direct and error-feedback compression operators, offering a thorough FedRL convergence analysis.
@@ 3 Preliminaries and Background
Discounted Infinite-horizon MDP.
@@
Optimal Policy and Bellman Operator.
@@
RL with a Generative Model.
@@ 5.1 CompFedRL with UnbiasedComp
Theorem 5.1 establishes non-asymptotic high-probability convergence with the linear rate depending on federated parameters $\beta,\eta,K$ and three residual error terms for CompFedRL with UnbiasedComp. We make the following observations. First, there is a trade-off between convergence speed and solution accuracy when we adjust the learning rate $\eta$. In particular, increasing $\eta$ improves the convergence speed while increasing the first residual error term. Second, the first and last error term depends on $1/\sqrt{I}$, thus implying that CompFedRL with UnbiasedComp attains the speed-up with respect to the number of agents like multi-agent TD(0) learning in 44. Third, the second error term is insignificant when the discount factor $\gamma$ is small. Fourth, Theorem 5.1 with $q_{2}=q_{\infty}=0$ provides the convergence for full-precision federated $Q$-learning.
@@ 6 Experiments
Setup.
We examine five variations of the noisy grid-world environment, some of which contain walls (as illustrated in Figure 5). Details regarding these environments and more results can be found in Appendix 0.C.1 and Appendix 0.C.2, respectively. The $Q$-function is initialized with zero entries, and we set $\gamma=0.8$. After each round of communication $t$, we evaluate the estimated $Q$ through the root mean square error (RMSE) between the estimated $\overline{Q}_{t}$ and the optimal $Q^{*}$.
@@
Communication Efficiency of CompFedRL.
@@
Impact of Federated Parameter $\beta$ and Learning Rate $\eta$.

### v2

@@ 1 Introduction
  • Empirical validation. To validate our theoretical observations, in Section 6, we conduct extensive numerical experiments, leveraging sparsification operators such as Top-$K$ and Sparsified-$K$. These experiments corroborate the effectiveness of CompFedRL in practice, further reinforcing the significance of our contributions in advancing communication-efficient FedRL algorithms. Our experiments’ code is publicly available in the following link: https://github.com/AliBeikmohammadi/CompFedRL.
@@ 2 Related Work
In what follows, we delve into the most pertinent threads of related work, with a particular focus on theoretical contributions in the field. We present prior works closely related to our paper in RL and communication-efficient learning.
2.1 Single-agent RL Algorithms
@@
2.2 Distributed and Federated RL Algorithms
@@
2.3 Communication-Efficient Learning Algorithms
Many compression schemes, i.e. sparsification and/or quantization, have been used for optimization algorithms to train state-of-the-art learning models (e.g. AlexNet, VGG19, and ResNet152) under resource-constrained networks [1, 67, 64, 44]. These compression schemes can be either unbiased (such as QSGD [1], TernGrad [67], and ATOMO [64]) or biased (e.g. Top-$K$[2, 7], normalized gradient [43] and sign compression [12]). Despite communication benefits, optimization algorithms with compression often have poor convergence performance. For instance, the biased compression algorithms are shown to fail to converge for simple problems [14, 36]. To address the problem of direct compression, several studies [54, 36, 2, 35, 41, 31, 46, 8] have proposed error-feedback mechanisms which leverage memory of previous compression errors to mitigate compression biases for distributed and federated learning algorithms under various settings. None of the mentioned works has investigated communication-efficient learning in the context of RL. The only existing study, [47], examines error-feedback compression, focusing on TD learning with linear function approximation in distributed learning, contrasting with our synchronous federated $Q$-learning approach, where aggregation occurs periodically. We investigate both direct and error-feedback compression operators, offering a thorough FedRL convergence analysis.
@@ 3 Preliminaries and Background
3.1 Discounted Infinite-horizon MDP
@@ 3.2 Policy, and $Q$-function
3.3 Optimal Policy and Bellman Operator
@@
3.4 RL with a Generative Model
@@ 5.1 CompFedRL with UnbiasedComp
Discussions on Theorem 5.1:
Theorem 5.1 establishes non-asymptotic high-probability convergence with the linear rate depending on federated parameters $\beta,\eta,K$ and three residual error terms for CompFedRL with UnbiasedComp. We make the following observations. First, there is a trade-off between convergence speed and solution accuracy when we adjust the learning rate $\eta$. In particular, increasing $\eta$ improves the convergence speed while increasing the first residual error term. Second, the first and last error term depends on $1/\sqrt{I}$, thus implying that CompFedRL with UnbiasedComp attains the speed-up with respect to the number of agents like multi-agent TD(0) learning in 47. Third, the second error term is insignificant when the discount factor $\gamma$ is small. Fourth, Theorem 5.1 with $q_{2}=q_{\infty}=0$ provides the convergence for full-precision federated $Q$-learning.
@@ 6 Experiments
6.1 Setup
We examine five variations of the noisy grid-world environment, some of which contain walls (as illustrated in Figure 5 of Appendix). Details regarding these environments and more results can be found in Appendix 0.C.1 and Appendix 0.C.2, respectively. The $Q$-function is initialized with zero entries, and we set $\gamma=0.8$. After each round of communication $t$, we evaluate the estimated $Q$ through the root mean square error (RMSE) between the estimated $\overline{Q}_{t}$ and the optimal $Q^{*}$.
@@ 6.1 Setup
6.2 Communication Efficiency of CompFedRL
@@ 6.3 Convergence Speedup
6.4 Impact of Federated Parameter $\beta$ and Learning Rate $\eta$
@@ 7 Conclusion
Acknowlegments
This work was partially supported by the Swedish Research Council through grant agreement no. 2020-03607 and in part by Sweden’s Innovation Agency (Vinnova). The computations were enabled by resources provided by the National Academic Infrastructure for Supercomputing in Sweden (NAISS) at Chalmers Centre for Computational Science and Engineering (C3SE) partially funded by the Swedish Research Council through grant agreement no. 2022-06725.
Supplementary Material
For detailed appendices and associated code, please refer to the full version of this paper available on Springer Link and arXiv [6], as well as the code repository on GitHub: https://github.com/AliBeikmohammadi/CompFedRL.